OpenStudy (anonymous):

find the equation of the line that passes through (1,-2) perpendicular to the line x+2y-1=0

OpenStudy (anonymous):

x + 2y - 1 = 0 y = (1 - x)/2 y' = -1/2 so the equation: y + 2 = (2)(x - 1) y = 2x - 4